Suppose that lo and I denote the light intensity before and after going through a material. Then according to Beer-Lambert's law"', x = -In () where k is a constant that depends on the material. If k = 0.03 and lo = 15 lumens, find the line intensity at a depth of 30 feet.


1
Expert's answer
2021-10-14T06:48:18-0400

According to Beer-Lambert’s law;

x=-(1/k)In(I/I0)

Given that I0=15 lumens

k=0.03

Light intensity at a depth of 30 feet

x=30ft

In(I/I0)=-k x

e In(I/I0)=e-k x

(I/IO)=e-k x

I(x)=I0e-kx

I(x)=I0e-0.03*10

=0.7408